OpenStudy (anonymous):

quadratic equation....need to show the steps x² - 2x – 13 = 0 and x² + 12x – 64 = 0

OpenStudy (bahrom7893):

First one: first find the discriminant: D = 4 - 4 * (-13) = 56 x1 = [ (-(-2) - sqrt(56) ] / 2

OpenStudy (amistre64):

x^2 + (b+c)x + (bc) = 0 is the format to use. b+c = -2 bc = -13 ; solve for b and c

OpenStudy (bahrom7893):

x2 = [ (-(-2) + sqrt(56) ] / 2

OpenStudy (amistre64):

it aint gonna play nice :) so complete the square maybe?

OpenStudy (bahrom7893):

Do the same steps for the next equation using the quadratics formula: ax^2 + bx + c = 0 D = b^2 - 4ac x1 = (-b - sqrt(D)) / (2a) x2 = (-b + sqrt(D)) / (2a)